OpenStudy (anonymous):

Subtract. –0.61 – (–7.74) A. –8.35 B. –7.13 C. 7.13 D. 8.35

OpenStudy (alexandervonhumboldt2):

hint: -0.61-(-7.74)=-0.61+7.74=7.74-0.61=?

OpenStudy (frozen15):

C. 7.13 because -(- changes to a + so its would be -0.61 + 7.74 = 7.13

OpenStudy (anonymous):

thanks

OpenStudy (frozen15):

np please medal :)

OpenStudy (frozen15):

fan me if you want and tag me in a post if any help needed for you or a friend , have a nice day!